OpenStudy (anonymous):

What is the resulting ordered pair if the value of the independent variable is -1/2 f(x)=12x^2-8x answer choose: A- [-1/2,-1] B- [-1/2,7] C- [-7,-1/2] D- [7,-1/2]

OpenStudy (anonymous):

@ganeshie8 please help

ganeshie8 (ganeshie8):

the independent variable is \(\large x\)

OpenStudy (anonymous):

yea?

ganeshie8 (ganeshie8):

plugin \(\large x = -1/2\) and evaluate the function

ganeshie8 (ganeshie8):

\(\large f(x) = 12x^2 - 8x\) \(\large f(-1/2) = ?\)

OpenStudy (anonymous):

[f(-1/2)=-3+4 \]

OpenStudy (anonymous):

right? @ganeshie8

ganeshie8 (ganeshie8):

you should get 3+4 because 12(-1/2)^2 = 12(1/4) = 3

ganeshie8 (ganeshie8):

so the y coordinate is 7

ganeshie8 (ganeshie8):

and the x coordinate is -1/2

ganeshie8 (ganeshie8):

an ordered pair is written as \(\large (x, y)\)

OpenStudy (anonymous):

so it would be B?

ganeshie8 (ganeshie8):

Yes!

OpenStudy (anonymous):

thank you so much! !!!!